Fermi National Laboratory


Accelerator Update - June 16-June 18

The integrated luminosity for last week was 5549.7 nb-1.

Monday June 16
The day shift began with Operations monitoring store 2689, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Operations terminated store 2689 at 12:08 PM and put the Tevatron (TeV) into shot setup fifty minutes later.

At 3 PM, Operations established store 2691 with an initial luminosity of 26.5E30. Pbar resumed stacking.

Booster experts tuned the Booster beam. This limited beam to MiniBooNE.

Equipment rack temperatures alarmed for a Pbar service building, AP-50. Operators installed some temporary fans at 6 PM.

Operators interlocked Switchyard enclosures F1, F2, and F3 at 7 PM, and then at 10:30 PM interlocked the external beamline enclosures M01, M02, M04, and M05. M03 had a problem that will be resolved by the safety system technicians in the morning.

Operators cooperated with a Main Injector (MI) expert in helping him conduct emittance studies from 10:30 to 11 PM.

Tuesday June 17
The midnight shift began with Operations monitoring store 2692, with Pbar stacking, and with MiniBooNE taking beam.

Operations saw several Linac quadrupole power supply trips. Operators had the problem fixed by 1:30 AM.

Operators turned down MiniBooNE at 1 AM and then began tuning on the Booster beam.

Operations terminated store 2692 at 6 AM and put the TeV into shot setup forty minutes later.

Two Booster RF stations (BRF12 & 14) tripped off at 7:34 AM. They both reset okay. Another station (BRF17) tripped of fifteen minutes later. It also reset okay.

Pbar suffered a separator spark at 8:30 AM.

At 8:36 AM, Operations established store 2694 with an initial luminosity of 29.98E30.

Operations turned off MI, Booster, and Linac at 8:47 AM and then made the areas ready for access. The MI access was made for Recycler work.

Operators interlocked external beamline enclosure M03.

The Linac access ended around 1 PM and the Booster access ended around 4 PM. MI was ready for beam by 5 PM.

The 400 MeV Debuncher, located in Linac but used by Booster, began causing problems at 5:05 PM due to high temperature. Experts investigated. There may have been a flow restriction that cleared itself. The LCW pump was also acting odd.

Operations had sent beam to the Recycler, Pbar, MI, and MiniBooNE by 5:49 PM.

Linac experts had the 400 MeV Debuncher working by 6:32 PM.

Operators were still tuning Booster at 10:30 PM. The problems were due to fluctuating RF cavity temperatures.

Wednesday June 18
The midnight shift began with Operations monitoring store 2694, with Pbar stacking, and with MiniBooNE taking beam.

The Tevatron Electron Lens (TEL) tripped off around 2 AM. Fortunately, Operations had warned the Collider Detector Facility and D-zero experiments that end-of-store studies were about to begin so the experiments had their high voltages off. A TEL expert had the device operating by 2:20 AM. The end-of-store studies then began.

Operations terminated store 2694 at 7:30 AM.

Operations put the TeV through a dry squeeze at 7:53 AM.

Today's Plan
The plan for today is to allow TeV experts to conduct beam maintenance studies until early afternoon. The experiments will be allowed an hour access after the studies end and then Operations will put the TeV into shot setup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/20/2003   email Fermilab